# Diffusion Models Meet Contextual Bandits

### v2

@@ Abstract
Efficient decision-making in contextual bandits with large action spaces is challenging, as methods lacking additional prior information may suffer from computational and statistical inefficiencies. In this work, we leverage pre-trained diffusion models as priors to capture complex action distributions and introduce a diffusion-based decision framework for contextual bandits. We develop practical algorithms to efficiently approximate posteriors under diffusion priors, enabling flexible decision-making strategies. Empirical evaluations demonstrate the effectiveness and versatility of our approach across diverse contextual bandit settings.
@@ 1 Introduction
A contextual bandit models online decision-making under uncertainty [49]. At each round, an agent observes a context, chooses an action, and receives a reward. The agent aims to maximize cumulative reward by balancing exploitation of high-reward actions and exploration of less-certain ones. The action space in contextual bandits is often large, resulting in less-than-optimal performance with standard exploration strategies (e.g., LinUCB [11] or LinTS [60]). Luckily, actions usually exhibit correlations, making efficient exploration possible as one action may inform the agent about other actions. In particular, Thompson sampling offers remarkable flexibility, allowing its integration with informative priors [33] that capture these correlations. Inspired by the achievements of diffusion models [62, 30], which effectively approximate complex distributions [23, 58], this work employs pre-trained diffusion models as priors in contextual Thompson sampling.
We illustrate the idea using video streaming. The objective is to optimize watch time for a user $j$ by selecting a video $a$ from a catalog of $K$ videos. Users $j$ and videos $a$ are associated with context vectors $x_{j}$ and unknown video parameters $\theta_{a}$, respectively. User $j$’s expected watch time for video $a$ is linear as $x_{j}^{\top}\theta_{a}$. Then, a natural strategy is to independently learn video parameters $\theta_{a}$ using LinTS or LinUCB [3, 1], but this proves statistically inefficient for larger $K$. Fortunately, the reward when recommending a movie can provide informative insights into other movies. To capture this, we leverage offline estimates of video parameters denoted by $\hat{\theta}_{a}$ and build a diffusion model on them. This pre-trained diffusion model approximates the video parameter distribution, capturing their dependencies. This model enriches contextual Thompson sampling as a prior, effectively capturing complex video dependencies while ensuring computational efficiency.
@@
We introduce a framework for contextual bandits with a diffusion-derived prior, and develop diffusion Thompson sampling (dTS) that is both computationally and statistically efficient. dTS achieves fast posterior updates and sampling through an efficient approximation that becomes exact when the diffusion prior and the likelihood are linear. A key contribution, beyond applying pre-trained diffusion models in contextual bandits, is the efficient computation and sampling of the posterior distribution of a $d$-dimensional parameter $\theta\mid\mathcal{D}$, with $\mathcal{D}$ representing the data, when using a pre-trained diffusion model prior on $\theta$. This is relevant not only to bandits and RL but also to a broader range of applications [19]. Our approximations are motivated with exact closed-form solutions obtained in cases where both the link functions of the pre-trained diffusion model and the likelihood are linear. These solutions form the basis for our approximations for non-linear link functions, demonstrating both strong empirical performance and computational efficiency. Our approach avoids the computational burden of heavy approximate sampling algorithms required for each latent parameter.
@@
Diffusion models have been applied to offline decision-making [6, 36, 65], but their use in online learning has only recently been explored by Hsieh et al. [34], who studied the multi-armed bandit setting, and by Kveton et al. [44], whose preprint appeared shortly after the first version of this work. An earlier and deliberately simplified version of our approach, corresponding to the fully linear-Gaussian case, was introduced in Aouali [7], which studied the simplest linear instance of diffusion-based contextual bandits and derived exact closed-form posteriors. That model laid the groundwork for the more general framework developed here. Our present study extends these efforts in two main directions. First, we generalize diffusion-based Thompson sampling to the more practical and realistic contextual bandit setting. Second, we show that when the diffusion model employs linear link functions and linear rewards, the resulting posteriors admit exact closed-form expressions without approximation. These exact posteriors are crucial: they enable theoretical analysis (unlike prior work) and motivate efficient approximations for non-linear link functions in contextual bandits. For an extended discussion of related work, see Appendix˜A.
@@ 2 Setting
Each action $a\in[K]$ is associated with an unknown action parameter $\theta_{a}\in^{d}$, so that the reward received in round $t$ is $Y_{t}\sim P(\cdot\mid X_{t};\theta_{A_{t}})$, where $P(\cdot\mid x;\theta_{a})$ is the reward distribution of action $a$ in context $x$. Throughout the paper, we assume that the reward distribution is parametrized as a generalized linear model (GLM) [54]. That is, $P(\cdot\mid x;\theta_{a})$ is an exponential-family distribution with mean $g(x^{\top}\theta_{a})$, where $g$ is the mean function. For example, we recover linear bandits when $P(\cdot\mid x;\theta_{a})=\mathcal{N}(\cdot;x^{\top}\theta_{a},\sigma^{2})$ where $\sigma>0$, and logistic bandits [24] with $g(u)=(1+\exp(-u))^{-1}$ and $P(\cdot\mid x;\theta_{a})=\mathrm{Ber}(g(x^{\top}\theta_{a}))$, where $\mathrm{Ber}(p)$ denotes the Bernoulli distribution with mean $p$.
@@
where $f_{\ell}$ and $\Sigma_{\ell}$ are known. Finally, the $L$-th latent parameter $\psi_{L}$ is sampled as
$\psi_{L}\sim\mathcal{N}(0,\Sigma_{L+1})\,$,
where $\Sigma_{L+1}$ is known. We summarize this model in Eq.˜1 below
@@ 3.1 Algorithm
$p(\theta_{a}\mid\psi_{1},H_{t,a})\propto\prod_{i\in S_{t,a}}P(Y_{i}\mid X_{i};\theta_{a})\,\mathcal{N}(\theta_{a};f_{1}(\psi_{1}),\Sigma_{1}),$
@@
$p(H_{t}\mid\psi_{1})=\prod_{a=1}^{K}\int_{\theta_{a}}\Bigg[\prod_{i\in S_{t,a}}P(Y_{i}\mid X_{i};\theta_{a})\Bigg]\mathcal{N}(\theta_{a};f_{1}(\psi_{1}),\Sigma_{1})\,\mathop{}\!\mathrm{d}\theta_{a},$
@@
$p(H_{t}\mid\psi_{\ell})=\int_{\psi_{\ell-1}}p(H_{t}\mid\psi_{\ell-1})\,\mathcal{N}(\psi_{\ell-1};f_{\ell}(\psi_{\ell}),\Sigma_{\ell})\,\mathop{}\!\mathrm{d}\psi_{\ell-1}.$
@@ 3.2 Posterior approximation
The reward distribution is parameterized as a generalized linear model (GLM) [54], allowing for non-linear rewards, which necessitates an approximation. We adopt an approach similar to the Laplace approximation, where a Gaussian density approximates the likelihood (instead of approximating the posterior). Specifically, the reward distribution $P(\cdot\mid x;\theta)$ belongs to the exponential family with mean function $g$. We approximate the likelihood as
$\prod_{i\in S_{t,a}}P(Y_{i}\mid X_{i};\theta_{a})\;\approx\;\mathcal{N}\!\big(\theta_{a};\hat{B}_{t,a},\hat{G}_{t,a}^{-1}\big),$
@@
$\hat{B}_{t,a}$ $=\operatorname*{arg\,max}_{\theta\in^{d}}\sum_{i\in S_{t,a}}\log P(Y_{i}\mid X_{i};\theta),\qquad\hat{G}_{t,a}=\sum_{i\in S_{t,a}}\dot{g}\!\big(X_{i}^{\top}\hat{B}_{t,a}\big)X_{i}X_{i}^{\top},$ (3)
and $S_{t,a}=\{\ell\in[t-1]:A_{\ell}=a\}$ is the set of rounds in which action $a$ was selected up to round $t$. Unlike Laplace, which approximates the entire posterior by a Gaussian, we only approximate the likelihood. This allows the approximate posterior to remain structurally a diffusion model, more expressive than a Gaussian, as described below.
After this initial approximation, we substitute the Gaussian likelihood into the posterior densities $p(\theta_{a}\mid\psi_{1},H_{t,a})$ and $p(\psi_{\ell-1}\mid\psi_{\ell},H_{t})$. This substitution removes the non-linearity of the reward model while retaining the non-linear diffusion links $f_{\ell}$. But, the resulting posteriors remain intractable when the link functions $f_{\ell}$ are non-linear. To address this, we introduce a second approximation inspired by the linear diffusion case, where $f_{\ell}(\psi_{\ell})=\mathrm{W}_{\ell}\psi_{\ell}$ with $\mathrm{W}_{\ell}\in^{d\times d}$ (see Section˜B.1). In that setting, closed-form posteriors can be derived (Section˜B.2), and we leverage these analytical solutions to construct efficient approximations by replacing the linear term $\mathrm{W}_{\ell}\psi_{\ell}$ with the general non-linear mapping $f_{\ell}(\psi_{\ell})$, yielding highly efficient posterior updates (see Appendix˜C).
Approximate action posterior. We approximate the conditional action posterior by
@@
$\hat{\Sigma}_{t,a}^{-1}$ $=\Sigma_{1}^{-1}+\hat{G}_{t,a},$ $\qquad\hat{\mu}_{t,a}$ $=\hat{\Sigma}_{t,a}\!\left(\Sigma_{1}^{-1}f_{1}(\psi_{1})+\hat{G}_{t,a}\hat{B}_{t,a}\right).$ (4)
In the absence of samples ($\hat{G}_{t,a}=0_{d\times d}$), the approximate posterior in Eq.˜4 exactly recovers the diffusion prior term $\mathcal{N}(f_{1}(\psi_{1}),\Sigma_{1})$ in Eq.˜1. As data accumulates, the likelihood term dominates, and the influence of the prior naturally diminishes.
Approximate latent posteriors. For each $\ell\in[L]\setminus\{1\}$, the $(\ell-1)$-th conditional latent posterior is approximated by
where
$\bar{\Sigma}_{t,\ell-1}^{-1}$ $=\Sigma_{\ell}^{-1}+\bar{G}_{t,\ell-1},$ $\qquad\bar{\mu}_{t,\ell-1}$ $=\bar{\Sigma}_{t,\ell-1}\!\left(\Sigma_{\ell}^{-1}f_{\ell}(\psi_{\ell})+\bar{B}_{t,\ell-1}\right).$ (5)
The top-layer latent posterior is
with
Recursive updates. The quantities $\bar{G}_{t,\ell}$ and $\bar{B}_{t,\ell}$ for $\ell\in[L]$ are computed recursively. The base recursion is
@@
With no data, $p(\psi_{\ell-1}\mid\psi_{\ell},H_{t})$ in Eq.˜5 coincides exactly with the diffusion prior $\mathcal{N}(f_{\ell}(\psi_{\ell}),\Sigma_{\ell})$ in Eq.˜1. As data accumulates, the prior’s influence diminishes. Thus, the approximation preserves a desirable property of exact posteriors: it matches the prior when no data is available, and the prior gradually fades as evidence grows.
Interpretation. This approximate posterior remains a diffusion model but with updated means and covariances. The latent-posterior means can be interpreted as updated link functions,
@@
and $\bar{\Sigma}_{t,\ell}$ as the corresponding updated covariances. Hence, this approximation yields a tractable yet expressive posterior (a diffusion model with updated parameters) without requiring heavy computations; unlike the Laplace approximation, which replaces the full posterior with a Gaussian. Alternative approaches such as variational inference would require solving complex optimization problems at each round, making them computationally prohibitive for bandit algorithms. Section˜G.3 provides empirical evidence that this approximation closely matches the exact posterior and compares favorably to these methods.
Remark 3.1 (Single shared action parameter).
The same algorithm and analysis extend to the case where all actions share a single unknown parameter $\theta\in^{d}$. Let $\varphi:\mathcal{X}\times[K]\to^{d}$ be a known feature map, and assume the reward model $g\!\left(\varphi(x,a)^{\top}\theta\right)$. Then, the diffusion prior in Eq.˜1 specializes by replacing the per-action parameters $(\theta_{a})_{a\in[K]}$ with a single $\theta$:
$\psi_{L}$ $\sim\mathcal{N}(0,\Sigma_{L+1}),$ (9)
A simple substitution $x^{\top}\theta_{a}\mapsto\varphi(x,a)^{\top}\theta$ and $(\theta_{a})_{a\in[K]}\mapsto\theta$ yield a single-parameter variant of dTS (theoretical results hold and similar approximation applies).

### v3

@@ Abstract
Efficient online decision-making in contextual bandits is challenging, as methods without informative priors often suffer from computational or statistical inefficiencies. In this work, we leverage pre-trained diffusion models as expressive priors to capture complex action dependencies and develop a practical algorithm that efficiently approximates posteriors under such priors, enabling both fast updates and sampling. Empirical results demonstrate the effectiveness and versatility of our approach across diverse contextual bandit settings.
@@ 1 Introduction
A contextual bandit models online decision-making under uncertainty [49]. At each round, an agent observes a context, selects an action, and receives a reward, aiming to maximize cumulative reward by balancing exploitation of high-reward actions and exploration of uncertain ones. However, in large-scale settings (e.g, the number of actions $K$ is large), standard exploration strategies (e.g., LinUCB [11] or LinTS [60]) often become computationally expensive or statistically inefficient. Fortunately, actions in many real-world problems exhibit correlations, enabling more efficient exploration since observing one action can inform the agent about others. Thompson sampling is particularly well-suited for this, as it naturally incorporates informative priors [33] that capture complex action dependencies. Inspired by the success of diffusion models [62, 30], which excel at approximating complex high-dimensional distributions [23, 58], this work leverages pre-trained diffusion models as priors in contextual Thompson sampling.
@@
Precisely, we introduce a framework for contextual bandits with a diffusion-derived prior, and develop diffusion Thompson sampling (dTS) that is both computationally and statistically efficient. dTS achieves fast posterior updates and sampling through an efficient approximation that becomes exact when the diffusion prior and the likelihood are linear. A key contribution, beyond applying pre-trained diffusion models in contextual bandits, is the efficient computation and sampling of the posterior distribution of a $d$-dimensional parameter $\theta\mid\mathcal{D}$, with $\mathcal{D}$ representing the data, when using a pre-trained diffusion model prior on $\theta$. This is relevant not only to bandits and RL but also to a broader range of applications [19]. Our approximations are motivated by exact closed-form solutions obtained in cases where both the pre-trained diffusion model and the likelihood are linear. These solutions form the basis for our approximations for the non-linear case, demonstrating both strong empirical performance and computational efficiency. Our approach avoids the computational burden of heavy approximate sampling algorithms required for each latent parameter.
@@
Diffusion models have been applied to offline decision-making [6, 36, 65], but their use in online learning has only recently been explored by Hsieh et al. [34] who studied the multi-armed bandit setting, and Kveton et al. [44] who explored a similar direction, with their preprint appearing shortly after the first version of this work. An earlier, deliberately simplified version of our approach, restricted to a linear diffusion model prior, was introduced in Aouali [7]. Here, we extend that formulation to the more realistic and expressive non-linear case. A detailed discussion of related work is provided in Appendix˜A.
@@ 2 Setting
We focus on the per-action (disjoint) setting, where each action $a\in[K]$ is represented by an unknown parameter vector $\theta_{a}\in\mathbb{R}^{d}$, so that the reward received in round $t$ is $Y_{t}\sim P(\cdot\mid X_{t};\theta_{A_{t}})$, where $P(\cdot\mid x;\theta_{a})$ is the reward distribution of action $a$ in context $x$. The reward distribution is parametrized as a generalized linear model (GLM) [54]. That is, $P(\cdot\mid x;\theta_{a})$ is an exponential-family distribution with mean $g(x^{\top}\theta_{a})$, where $g$ is the mean function. For example, we recover linear bandits when $P(\cdot\mid x;\theta_{a})=\mathcal{N}(\cdot;x^{\top}\theta_{a},\sigma^{2})$ where $\sigma>0$, and logistic bandits [24] with $g(u)=(1+\exp(-u))^{-1}$ and $P(\cdot\mid x;\theta_{a})=\mathrm{Ber}(g(x^{\top}\theta_{a}))$, where $\mathrm{Ber}(p)$ denotes the Bernoulli distribution with mean $p$. All derivations and algorithms extend naturally to the shared-parameter case described in Remark˜2.1.
@@
where $f_{\ell}$ and $\Sigma_{\ell}$ are known. Finally, the $L$-th latent parameter $\psi_{L}$ is sampled as $\psi_{L}\sim\mathcal{N}(0,\Sigma_{L+1})$, where $\Sigma_{L+1}$ is known. We summarize this model in Eq.˜1 below
@@
Remark 2.1 (Single shared action parameter).
Our algorithm and analysis also apply to the case where all actions share a single unknown parameter $\theta\in^{d}$. Let $\varphi:\mathcal{X}\times[K]\to^{d}$ be a known feature map, and assume the reward distribution mean is $g\!\left(\varphi(x,a)^{\top}\theta\right)$. Then, the diffusion prior in Eq.˜1 specializes by replacing the per-action parameters $(\theta_{a})_{a\in[K]}$ with a single shared parameter $\theta$:
$\psi_{L}$ $\sim\mathcal{N}(0,\Sigma_{L+1}),$ (2)
This formulation is useful when a shared feature map $\varphi$ is available. In that case, the diffusion model can be pre-trained on parameters $\{\theta_{s}\}_{s=1}^{S}$ from previous tasks, and dTS can then be applied to a new task S+1S{+}1 using the pre-trained prior. To avoid clutter, our main exposition focuses on the model in Eq.˜1, but all theoretical results and algorithmic components extend naturally to this shared-parameter case, which we also include in some experiments (explicitly noted when applicable).
@@ 3.1 Algorithm
$p(\theta_{a}\mid\psi_{1},H_{t,a})\propto\prod_{i\in S_{t,a}}P(Y_{i}\mid X_{i};\theta_{a})\,\mathcal{N}(\theta_{a};f_{1}(\psi_{1}),\Sigma_{1}),$ (4)
@@
$p(H_{t}\mid\psi_{1})=\prod_{a=1}^{K}\int_{\theta_{a}}\Bigg[\prod_{i\in S_{t,a}}P(Y_{i}\mid X_{i};\theta_{a})\Bigg]\mathcal{N}(\theta_{a};f_{1}(\psi_{1}),\Sigma_{1})\,\mathop{}\!\mathrm{d}\theta_{a},$ (5)
@@
$p(H_{t}\mid\psi_{\ell})=\int_{\psi_{\ell-1}}p(H_{t}\mid\psi_{\ell-1})\,\mathcal{N}(\psi_{\ell-1};f_{\ell}(\psi_{\ell}),\Sigma_{\ell})\,\mathop{}\!\mathrm{d}\psi_{\ell-1}.$ (6)
@@ 3.2 Posterior approximation
The reward distribution is parameterized as a generalized linear model (GLM) [54], which allows for non-linear rewards. In addition, the diffusion model itself is highly non-linear due to the link functions $f_{\ell}$. These two sources of non-linearity make the posterior intractable, so we apply two layers of approximation: (i) a likelihood approximation to linearize the reward model, and (ii) a diffusion approximation to handle the non-linear hierarchy induced by the diffusion model prior.
(i) Likelihood approximation. We use an approach similar to the Laplace approximation, but instead of approximating the entire posterior, we approximate only the likelihood by a Gaussian. Precisely, the reward distribution $P(\cdot\mid x;\theta_{a})$ belongs to the exponential family with mean function $g$. Thus
$\prod_{i\in S_{t,a}}P(Y_{i}\mid X_{i};\theta_{a})\;\approx\;\mathcal{N}\!\big(\theta_{a};\hat{B}_{t,a},\hat{G}_{t,a}^{-1}\big),$ (7)
@@
$\hat{B}_{t,a}$ $=\operatorname*{arg\,max}_{\theta_{a}\in^{d}}\sum_{i\in S_{t,a}}\log P(Y_{i}\mid X_{i};\theta_{a}),\qquad\hat{G}_{t,a}=\sum_{i\in S_{t,a}}\dot{g}\!\big(X_{i}^{\top}\hat{B}_{t,a}\big)X_{i}X_{i}^{\top},$ (8)
and $S_{t,a}=\{\ell\in[t-1]:A_{\ell}=a\}$ is the set of rounds in which action $a$ was selected. Unlike Laplace, which fits a global Gaussian to the full posterior, this step only linearizes the local likelihood, allowing the hierarchical diffusion structure of the prior to remain intact and expressive.
(ii) Diffusion approximation. Plugging the Gaussian likelihood approximation (7) into the posterior expressions $p(\theta_{a}\mid\psi_{1},H_{t,a})$ and $p(\psi_{\ell-1}\mid\psi_{\ell},H_{t})$ removes the non-linearity of the reward model. However, the diffusion hierarchy remains non-linear through $f_{\ell}$. To handle this, we build on the closed-form posteriors of the linear diffusion case (where $f_{\ell}(\psi_{\ell})=\mathrm{W}_{\ell}\psi_{\ell}$; see Appendix˜B) and generalize them by replacing the linear terms $\mathrm{W}_{\ell}\psi_{\ell}$ with their non-linear counterparts $f_{\ell}(\psi_{\ell})$. This substitution yields a posterior diffusion model that retains the same hierarchical form as the prior but with data-dependent means and covariances. Details on how we transition from the linear to the general non-linear setting are provided in Appendices˜B and C. The resulting approximate posteriors (both action and latent) admit the following closed-form expressions.
Approximate action posterior. We approximate the conditional action posterior as
@@
where
$\hat{\Sigma}_{t,a}^{-1}$ $=\underbrace{\Sigma_{1}^{-1}}_{\text{prior precision}}\;+\;\underbrace{\hat{G}_{t,a}}_{\text{data precision}},$ $\qquad\hat{\mu}_{t,a}$ $=\hat{\Sigma}_{t,a}\!\Big(\underbrace{\Sigma_{1}^{-1}f_{1}(\psi_{1})}_{\text{prior contribution}}\;+\;\underbrace{\hat{G}_{t,a}\hat{B}_{t,a}}_{\text{data contribution}}\Big).$ (9)
This posterior update has a clear interpretation. The posterior precision $\hat{\Sigma}_{t,a}^{-1}$ is the sum of the prior precision and the data precision. The posterior mean $\hat{\mu}_{t,a}$ is the precision-weighted average of the prior mean and the MLE $\hat{B}_{t,a}$. As more data are observed, the covariance shrinks and the mean moves from the prior mean $f_{1}(\psi_{1})$ toward the MLE $\hat{B}_{t,a}$. When no data are available ($\hat{G}_{t,a}=0$), the posterior reduces to the prior $\mathcal{N}(f_{1}(\psi_{1}),\Sigma_{1})$; in the limit of infinite data ($\hat{G}_{t,a}\!\to\!\infty$), the posterior collapses to the MLE $\hat{B}_{t,a}$, with $\hat{\mu}_{t,a}\!\to\!\hat{B}_{t,a}$ and $\hat{\Sigma}_{t,a}\!\to\!0$.
Approximate latent posteriors. For each $\ell\in[L+1]\setminus\{1\}$, we approximate the latent posterior as
@@
$\bar{\Sigma}_{t,\ell-1}^{-1}=\underbrace{\Sigma_{\ell}^{-1}}_{\text{prior precision}}+\underbrace{\bar{G}_{t,\ell-1}}_{\text{data precision}},\;\bar{\mu}_{t,\ell-1}=\bar{\Sigma}_{t,\ell-1}\!\Big(\underbrace{\Sigma_{\ell}^{-1}f_{\ell}(\psi_{\ell})}_{\text{prior contribution}}+\underbrace{\bar{B}_{t,\ell-1}}_{\text{data contribution}}\Big)\,$, (10)
where, by convention, $f_{L+1}(\psi_{L+1})=0$ since the top layer $\psi_{L}$ has no parent $\psi_{L+1}$. The quantities $\bar{G}_{t,\ell}$ and $\bar{B}_{t,\ell}$ are computed recursively. The base recursion is
@@
The latent posterior update in Eq.˜10 has the same structure as the action posterior. The posterior precision $\bar{\Sigma}_{t,\ell-1}^{-1}$ is the sum of the prior and data precisions , and the posterior mean is their precision-weighted combination. The data terms $\bar{G}_{t,\ell-1}$ and $\bar{B}_{t,\ell-1}$ are computed recursively (Eqs.˜11 and 12), so information collected at the action level propagates upward through the hierarchy.
Interpretation. The resulting approximate posterior remains a diffusion model whose conditional Gaussians have updated, data-dependent means and covariances. The latent-posterior means can be viewed as refined link functions:
@@
and $\bar{\Sigma}_{t,\ell}$ represents their updated uncertainty. Both are updated with data: covariances contract as uncertainty decreases, and means move from the prior toward the MLE. Unlike a full Laplace approximation, this formulation preserves the expressiveness of the posterior rather than replacing it globally with a single Gaussian, while also avoiding the heavy computation required by other approximate inference methods.
3.3 Extension to single shared action parameter
For the shared-parameter model in Remark˜2.1, dTS’s posterior approximations are similar. The action posterior is $p(\theta\mid\psi_{1},H_{t})\approx\mathcal{N}(\hat{\mu}_{t},\hat{\Sigma}_{t})$, where
$\hat{\Sigma}_{t}^{-1}=\Sigma_{1}^{-1}+\hat{G}_{t}\,$, $\hat{\mu}_{t}=\hat{\Sigma}_{t}\!\big(\Sigma_{1}^{-1}f_{1}(\psi_{1})+\hat{G}_{t}\hat{B}_{t}\big).$ (13)
where
$\hat{B}_{t}=\operatorname*{arg\,max}_{\theta\in\mathbb{R}^{d}}\sum_{i<t}\log P\!\left(Y_{i}\mid\varphi(X_{i},A_{i})^{\top}\theta\right),$ $\hat{G}_{t}=\sum_{i<t}\dot{g}\!\big(\varphi(X_{i},A_{i})^{\top}\hat{B}_{t}\big)\,\varphi(X_{i},A_{i})\varphi(X_{i},A_{i})^{\top}.$
Similarly, for $\ell\in[L+1]\setminus\{1\}$, the latent posterior is $p(\psi_{\ell-1}\mid\psi_{\ell},H_{t})\approx\mathcal{N}(\bar{\mu}_{t,\ell-1},\bar{\Sigma}_{t,\ell-1})$, where
$\bar{\Sigma}_{t,\ell-1}^{-1}$ $=\Sigma_{\ell}^{-1}+\bar{G}_{t,\ell-1},\qquad\bar{\mu}_{t,\ell-1}=\bar{\Sigma}_{t,\ell-1}\!\big(\Sigma_{\ell}^{-1}f_{\ell}(\psi_{\ell})+\bar{B}_{t,\ell-1}\big)\,$, (14)
where, by convention, $f_{L+1}(\psi_{L+1})=0$ and the quantities $\bar{G}_{t,\ell}$ and $\bar{B}_{t,\ell}$ are computed recursively as
Base case: $\bar{G}_{t,1}=\Sigma_{1}^{-1}-\Sigma_{1}^{-1}\hat{\Sigma}_{t}\Sigma_{1}^{-1},\qquad\bar{B}_{t,1}=\Sigma_{1}^{-1}\hat{\Sigma}_{t}\hat{G}_{t}\hat{B}_{t}.$ (15)
Recursive case: $\bar{G}_{t,\ell}=\Sigma_{\ell}^{-1}-\Sigma_{\ell}^{-1}\bar{\Sigma}_{t,\ell-1}\Sigma_{\ell}^{-1},\qquad\bar{B}_{t,\ell}=\Sigma_{\ell}^{-1}\bar{\Sigma}_{t,\ell-1}\bar{B}_{t,\ell-1}.$ (16)
Again, this shared-parameter variant of dTS is presented for completeness and to illustrate the generality of our posterior derivations; the main focus of the paper remains on the per-action formulation in Eq.˜1. Unless stated otherwise, all theoretical results and experiments use the main version of dTS described in Algorithm˜1.